OpenStudy (anonymous):

PLEASE HELP!!! What is the simplified form of y^2/x over y^5/x^2

OpenStudy (anonymous):

\[\frac{ y ^{2} }{ x } \over \frac{ y ^{5} }{ x ^{2} }\]

OpenStudy (anonymous):

a) y^7/x^3 b) 1/xy^3 c) xy3 d) x/y^3

OpenStudy (anonymous):

use this: \(\large \frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b} \cdot \frac{d}{c} \) then simplify the problem..

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Its like 4/ (1/4) which is just like 4/.25 which if you put in a claculator you get 16, well in frction form we get \[\Huge \frac{\frac{4}{1}}{\frac{1}{4}}\]

OpenStudy (anonymous):

Well its the same thing as doing this if we let a-4 b-1 c-1 d-1 \[\Large \frac{4}{1} \times \frac{4}{1}\]

OpenStudy (anonymous):

Are you sure? Can you show me what you did? Like in the fraction form?

OpenStudy (psymon):

Again, separate the two fractions and have them side by side in a way you're more accustomed to visually looking at. You would have (y^2/x) divided by (y^5/x^2) when you divide fractions, you'll take the second fraction and flip it then multiply across. So this problem turns into (y^2/x) * (x^2/y^5). This leads to (x^2y^2/xy^5) Simplifying this down leaves you with (x/y^3)